OpenStudy (anonymous):

find the sum of the first 20 terms in the arithmetic series -6-12-18

OpenStudy (anonymous):

If a series is arithmetic the sum of the first n terms, denoted Sn, there are ways to find its sum without actually adding all of the terms. To find the sum of the first n terms of an arithmetic series use the formula, , where n is the number of terms, a1 is the first term and an is the last term. Example 1: Find the sum of the first 20 terms of the arithmetic series if a1 = 5 and a20 = 62. Example 2: Find the sum of the first 40 terms of the arithmetic series 2 + 5 + 8 + 11 + ···. First find the 40th term: Then find the sum: Example 3: Find the sum: First find a1 and a50: a1 = 3(1) + 2 = 5 a50 = 3(50) + 2 = 152 Then find the sum:

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

yea kind of

OpenStudy (anonymous):

What else do you need?

geerky42 (geerky42):

Nice copy/paste @bleebleebleed http://hotmath.com/hotmath_help/topics/sum-of-the-first-n-terms-of-an-arithmetic-series.html

OpenStudy (anonymous):

I know i did

OpenStudy (anonymous):

At least i took the time to do it @geerky42

geerky42 (geerky42):

Just so you can get medal, right...? @cutegirl To find sum , you use this formula: \(S_n = \dfrac{n(a_n+a_1)}{2}\) So all you need is first term (-6) and 20th term. You have sequence -6,-12,-18,... Can you find 20th term?

OpenStudy (anonymous):

Nope I didnt want a medal